Grade 11Engineering Entrance Exams

If the volume of block of metal changes by 0.12% when it is heated through 20°c what is it's coefficient of linear expansion_OC`

Akhil
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
 
Dear Akhil
 
Y= dV /V * dT
 
now you will get Y (gamma) = 6 * 10^-5
 
hence coefficient of linear expansion is = Y/3 = 2 * 10^-5 degree celsius

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago